OpenStudy (anonymous):

10. (LEVEL 3)Simplify 11. (LEVEL 3)Robert loves football and wants to start practice throwing daily. The first day he practices for 3 minutes. The next day he triples his time. Robert continues to triple his practice time every day. How many minutes did Robert practice on the fourth day? 3x3x3= 27 27x27x27=1683 12. (LEVEL 3)Given a cube whose faces each have area A, the volume of the cube is given by the formula V = A3/2. Find the volume of a cube whose faces each have an area of 16 in2. 13. (LEVEL 4) Simplify 14. (LEVEL 4) Simplify

OpenStudy (anonymous):

@supercrazy92 @samtasticc @SWAG @henpen @Hope99 @Hayden14 @Jonask @jen1511 @Jenn777 @jasmynarteaga1 @samtasticc

OpenStudy (anonymous):

Dear friend... I'm terribly bad at math :P But good luck!

OpenStudy (anonymous):

copy and paste error wat happened to 10,13,14

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

@SamuelAlden917 @seiga @saifoo.khan @amistre64 @Agent_Sniffles

OpenStudy (anonymous):

@Winnie.Leo97 @walters @whpalmer4 @qaisar900 @experimentX @lgg23 @pinkpearl12000 @Piggiepigpigs

OpenStudy (anonymous):

@Winnie.Leo97 @walters @whpalmer4 @experimentX @experimentX @emma_beiber @ryan123345 @Reppinthe80 @rajathsbhat @redred413 @calmat01 @CoughCoughAchoo @seiga @Chooch146 @Chooch146

OpenStudy (anonymous):

Hmmm...perhaps you are meant to do this on your own since it is an exit assessment. All I see are a bunch of x's where your pictures should be.

OpenStudy (anonymous):

i just need help and to check my anwsers?

OpenStudy (experimentx):

3+3x3+3x3x3+3x3x3x3 for 1

OpenStudy (experimentx):

for the next \[ (16)^{3/2} = 4^3 = 64\]

OpenStudy (anonymous):

is that for the cube one??

OpenStudy (experimentx):

yeah ... sorry a bit busy for a while.

OpenStudy (anonymous):

its okay